I have tried every way to solve this and I cannot get it to work. Please help.
3(x-7) = 2x - 11
I tried 3x - 21 = 2x - 11
which becomes 3x- 2x - 21 = 2x - 2x -11
therefore 1x - 21 = 11
cant get it to work so I tried 5x -21 = 11 still didnt make sense
V_SNOOKS@sky.com

Answer by Alan3354(69443)

3(x-7) = 2x - 11
I tried 3x - 21 = 2x - 11
which becomes 3x- 2x - 21 = 2x - 2x -11
therefore 1x - 21 = 11 **** = -11
x - 21 = -11
Add 21
x = 10